# CrossLMM: Decoupling Long Video Sequences from LMMs via Dual Cross-Attention Mechanisms

### v1

@@ 4.2 Comparison with State-of-the-art Methods
In this section, we present a comparative analysis of CrossLMM against state-of-the-art LMMs, encompassing commercial implementations [41, 42, 44], open-source LMMs [9, 29, 59, 72, 7, 6, 56, 78, 26, 79], and specialized open-source long video LMMs [31, 35, 62, 73, 57, 46].
@@
[…]
Model Size #tokens MVBench PerceptionTest LongVideoBench MLVU VideoMME VideoMME per frame Avg Val Val M-Avg w/o sub. w sub. Avg. Duration 16s 23s 473s 651s 1010s 1010s Proprietary Models GPT4-V [41] - - 43.7 - 59.1 49.2 59.9 63.3 GPT4-o [42] - - 64.6 - 66.7 64.6 71.9 77.2 Gemini-1.5-Pro [44] - - 60.5 - 64.0 - 75.0 81.3 Small Size LMMs Qwen2-VL [56] 2B - 63.2 - - - 55.6 60.4 InternVL2.5 [6] 2B 256 68.8 - 46.0 61.4 51.9 54.1 CrossLMM 2B 1 58.6 58.1 47.3 57.6 55.5 58.9 CrossLMM 2B 9 63.5 63.5 51.8 61.0 59.1 61.3 Open-Source LMMs VideoLLaMA2 [9] 7B 72 54.6 51.4 - 48.5 47.9 50.3 VideoLLaMA2 [9] 72B 72 62.0 57.5 - - 62.4 64.7 VideoChat2-HD [29] 7B 72 62.3 - - 47.9 45.3 55.7 InternVideo2-HD [59] 7B 72 67.2 63.4 - - 49.4 - IXComposer-2.5 [72] 7B 400 69.1 34.4 - 37.3 55.8 58.8 InternVL2 [7] 8B 256 65.8 - 54.6 64.0 54.0 56.9 InternVL2 [7] 76B 256 69.6 - 61.1 69.9 61.2 62.8 InternVL2.5 [6] 8B 256 72.0 - 60.0 68.9 64.2 66.9 Qwen2-VL [56] 7B - 67.0 62.3 - - 63.3 69.0 Qwen2-VL [56] 72B - 73.6 68.0 - - 71.2 77.8 LLaVA-NeXT-Video [78] 7B 144 53.1 48.8 49.1 - - 46.5 LLaVA-OneVision [26] 7B 196 56.7 57.1 56.3 64.7 58.2 61.5 LLaVA-OneVision [26] 72B 196 59.4 66.9 61.3 68.0 66.2 69.5 LLaVA-Video [79] 7B 676 58.6 67.9 58.2 70.8 63.3 69.7 Open-Source Long Video LMMs LLaMA-VID [31] 7B 2 41.9 44.6 - 33.2 25.9 - Kangaroo [35] 8B 256 61.0 - 54.8 61.0 56.0 57.6 LongVILA [62] 7B 196 67.1 58.1 57.1 - 60.1 65.6 LongVA [73] 7B 144 - - - 56.3 52.6 54.3 LongLLaVA [57] 9B 144 49.1 - - - 43.7 - LongVU [46] 7B 64 66.9 - - 65.4 - 60.6 CrossLMM 7B 1 62.7 64.8 54.5 64.0 61.3 63.7 CrossLMM 7B 9 68.2 68.4 56.0 67.2 62.6 64.7 CrossLMM  7B 16 68.8 71.4 60.4 70.7 65.4 67.6
@@ 4.3 Efficiency Analysis
Figure 4 presents a comprehensive efficiency comparison between LLaVA-OV [26] and CrossLMM across different frame numbers (32, 64, 128, and 256) with 8 H800. The analysis focuses on three critical metrics: CUDA memory consumption, computational complexity, and prefill time.
@@
Processing Time Efficiency. The prefill processing time, depicted in Figure 4(c), shows that CrossLMM consistently outperforms LLaVA-OV [26] in terms of speed. At 32 frames, CrossLMM completes processing in 335ms, which is 39.6% faster than LLaVA-OV’s 554.91ms. This advantage is maintained across higher frame counts, with CrossLMM processing 256 frames in 1,975ms compared to LLaVA-OV’s 3,978.83ms, representing a 50.4% reduction in processing time. The time efficiency of CrossLMM is particularly significant for real-time applications where latency is a critical factor.
4.4 Ablation Study
@@ 5 Conclusion
We present CrossLMM, which offers a rigorous solution to the computational inefficiency that has long constrained LMMs when processing extended video sequences. Our CrossLMM framework demonstrates that through judicious token reduction strategies and architectural innovations, significant computational efficiency can be achieved while preserving model fidelity. The dual cross-attention mechanism, comprising V2V cross attention for maintaining fine-grained representational integrity and T2V cross attention for enhanced multimodal comprehension, constitutes a methodological advancement in token management for the practical real-world multimodal systems.
Appendix A Implementation Details
A.1 Training Strategies and Data
@@
In contrast to current mainstream LMMs [56, 6, 7], which typically employ intricate three- or four-stage training pipelines, we argue that such methodologies introduce unnecessary complexity and accessibility barriers. To address this, we adopt a streamlined two-stage training framework inspired by earlier LMM practices [34, 83].
@@
  • Stage I: Pretraining Data. The pre-training corpus for CrossLMM consists of two primary components: (1) 15M image-text pairs from CapsFusion [70], and (2) 5M video-text pairs sourced from InternVid [58]. A sampling ratio of 1:1 is implemented between image-text and video-text pairs throughout the training process.
@@
    • Image Instruction data. In our methodology, we employed a corpus of single-image instruction data derived from multiple established datasets, specifically LLaVA-NeXT [78], ALLaVA [5], and ShareGPT4-o [59, 7]. To enhance the model’s capacity for processing complex visual inputs, we further augmented our training regime with multi-image sequential data obtained from LLaVA-Interleave [27]. This comprehensive data integration approach facilitated a more robust visual-linguistic representational framework.
    • Short Video Instruction data. For the instruction fine-tuning phase of our research, we predominantly employed short video sequences from VideoChat2 [29] and InternVideo2 [59] datasets. To enhance the model’s comprehension capabilities, we further supplemented the training corpus with annotations derived via GPT4-o from several established datasets, including ShareGPT4o [59, 7], VideoChatGPT-Plus  [40], LLaVA-Video-178K [79], and LLava-Hound [76]. This methodical integration of diverse video-based instructional data facilitated the development of a more comprehensive temporal-visual understanding framework.
    • Long Video Instruction data. In our experimental framework, we primarily utilized long-form video instruction datasets, specifically those sourced from MovieChat [49] and Vript [67], supplemented by LongVid corpus [30]. This methodological approach to data selection enabled comprehensive training on extended temporal sequences, thereby facilitating the model’s capacity to process, interpret, and generate responses to complex narrative structures and sustained visual information across prolonged video segments.
A.2 Training hyperparameters.
@@
Vision Resolution×$\times$×Num. frames 384 384 $\times$8
#Tokens 9 $\times$ 32 9 $\times$ (64$\sim$512)
@@
LR of vision encoder 1×$\times 10^{-3}$ 2×$\times 10^{-6}$
LR of connector & LLM 1×$\times 10^{-3}$ 1 ×$\times 10^{-5}$
@@
Appendix B More Experiments
B.1 Ablation Study.
Pretraining data volume analysis.
@@ Appendix B More Experiments
B.2 Computational Efficiency Analysis
@@
Appendix C Limitation and Future Work

### v2

@@ 4.2 Comparison with State-of-the-art Methods
In this section, we present a comparative analysis of CrossLMM against state-of-the-art LMMs, encompassing commercial implementations [gpt4v, gpt4o, gemini], open-source LMMs [videollama2, videochat2, internvideo2, xcomposer, internvl2, internvl2_5, qwen2vl, llavanextvideo, llavaonevision, llavavideo], and specialized open-source long video LMMs [llamavid, kangaroo, longvila, longva, longllava, longvu].
@@ 4.3 Efficiency Analysis
Figure 4 presents a comprehensive efficiency comparison between LLaVA-OV [llavaonevision] and CrossLMM across different frame numbers (32, 64, 128, and 256) with 8 H800. The analysis focuses on three critical metrics: CUDA memory consumption, computational complexity, and prefill time.
@@
Processing Time Efficiency. The prefill processing time, depicted in Figure 4(c), shows that CrossLMM consistently outperforms LLaVA-OV [llavaonevision] in terms of speed. At 32 frames, CrossLMM completes processing in 335ms, which is 39.6% faster than LLaVA-OV’s 554.91ms. This advantage is maintained across higher frame counts, with CrossLMM processing 256 frames in 1,975ms compared to LLaVA-OV’s 3,978.83ms, representing a 50.4% reduction in processing time. The time efficiency of CrossLMM is particularly significant for real-time applications where latency is a critical factor.
4.4 Ablation Study
@@ 5 Conclusion
We present CrossLMM, which offers a rigorous solution to the computational inefficiency that has long constrained LMMs when processing extended video sequences. Our CrossLMM framework demonstrates that through judicious token reduction strategies and architectural innovations, significant computational efficiency can be achieved while preserving model fidelity. The dual cross-attention mechanism, comprising visual-to-visual cross attention for maintaining fine-grained representational integrity and text-to-visual cross attention for enhanced multimodal comprehension, constitutes a methodological advancement in token management for practical real-world multimodal systems.
[…]
Supplementary Material
6 Implementation Details
6.1 Training Strategies and Data
@@
In contrast to current mainstream LMMs [qwen2vl, internvl2_5, internvl2], which typically employ intricate three- or four-stage training pipelines, we argue that such methodologies introduce unnecessary complexity and accessibility barriers. To address this, we adopt a streamlined two-stage training framework inspired by earlier LMM practices [llava, minigpt4].
@@
  • Stage I: Pretraining Data. The pre-training corpus for CrossLMM consists of two primary components: (1) 15M image-text pairs from CapsFusion [capsfusion], and (2) 5M video-text pairs sourced from InternVid [internvid]. A sampling ratio of 1:1 is implemented between image-text and video-text pairs throughout the training process.
@@
    • Image Instruction data. In our methodology, we employed a corpus of single-image instruction data derived from multiple established datasets, specifically LLaVA-NeXT [llavanextvideo], ALLaVA [allava], and ShareGPT4-o [internvideo2, internvl2]. To enhance the model’s capacity for processing complex visual inputs, we further augmented our training regime with multi-image sequential data obtained from LLaVA-Interleave [llavainterleave]. This comprehensive data integration approach facilitated a more robust visual-linguistic representational framework.
    • Short Video Instruction data. For the instruction fine-tuning phase of our research, we predominantly employed short video sequences from VideoChat2 [videochat2] and InternVideo2 [internvideo2] datasets. To enhance the model’s comprehension capabilities, we further supplemented the training corpus with annotations derived via GPT4-o from several established datasets, including ShareGPT4o [internvideo2, internvl2], VideoChatGPT-Plus  [videogpt+], LLaVA-Video-178K [llavavideo], and LLava-Hound [llavahound]. This methodical integration of diverse video-based instructional data facilitated the development of a more comprehensive temporal-visual understanding framework.
    • Long Video Instruction data. In our experimental framework, we primarily utilized long-form video instruction datasets, specifically those sourced from MovieChat [moviechat] and Vript [vript], supplemented by LongVid corpus [li2024videochat]. This methodological approach to data selection enabled comprehensive training on extended temporal sequences, thereby facilitating the model’s capacity to process, interpret, and generate responses to complex narrative structures and sustained visual information across prolonged video segments.
6.2 Training hyperparameters.
@@
[…]
@@
[…]
LR of connector & LLM $\times 10^{-3} 1$ $\times 10^{-5}$
@@
7 More Experiments
7.1 Ablation Study.
Pretraining data volume analysis.
Vision encoder trainability investigation.
7.2 Computational Efficiency Analysis
@@
8 Limitation and Future Work